OpenStudy (anonymous):

HELP PLZ how many real solutions does the equation have -8x^2-8x-2=0

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

shall we find them? then we will know for sure !

OpenStudy (anonymous):

A: one solution B:two solutions c:no solutions D: infinite

OpenStudy (misty1212):

i would divide all by \(-2\) so make life easier and start with \[4x^2+4x-1=0\] then compute the discriminant or just use the quadratic formula which has the discriminant in it

OpenStudy (misty1212):

also i made a mistake, it should be \[\huge 4x^2+4x+1=0\]

OpenStudy (misty1212):

you might even recognize this as a perfect square, but even if you do not, you can compute \[b^2-4ac\] with \[a=4,b=4,c=1\]

OpenStudy (misty1212):

let me know what you get

OpenStudy (anonymous):

ok let me do the math real quick

OpenStudy (misty1212):

yeah good "do the math" is pretty much what this is all about

OpenStudy (anonymous):

i got 4+- sqrt of 0/8

OpenStudy (misty1212):

lol

OpenStudy (anonymous):

?...

OpenStudy (misty1212):

\[b^2-4ac=4^2-16=0\] so there is one solution

OpenStudy (anonymous):

oh ok i get it i put it over 4a cuz thats the quadratic formula

OpenStudy (misty1212):

do not write \(\sqrt0\) your math teacher will think you are a moron, just leave it as \[\frac{4}{8}=\frac{1}{2}\]

OpenStudy (tkhunny):

-8x^2-8x-2=0 Dividing by -2 4x^2 + 4x + 1 = 0 Immediately, there are NO positive Real Roots Substituting -x 4x^2 - 4x + 1 = 0 There might be 2 negative Real Roots (Counting multiplicities) Discriminant \(4^{2} - 4(4)1 = 16 - 16 = 0\) Aha!! There are two negative roots and they are equal. Sometimes, the spirit of the question does not require the solution. Just answer the question.

OpenStudy (misty1212):

in fact \[4x^2+4x+1=(2x+1)^2\] and if \[(2x+1)^2=0\] then the only solution is \[2x+1=0\\ 2x=-1\\ x=-\frac{1}{2}\]

OpenStudy (anonymous):

so then there is 1 solution..

OpenStudy (misty1212):

yes dear just one be careful here

OpenStudy (anonymous):

ok thank you :) could u help me with another?

OpenStudy (misty1212):

the quadratic formula is \[\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]you dropped the minus sign

OpenStudy (misty1212):

yes sure why not?

OpenStudy (anonymous):

what is the value of C so that x^2-11x+c is a perfect-square trinomial

OpenStudy (misty1212):

what is half of 11?

OpenStudy (misty1212):

use a fraction, not a decimal

OpenStudy (tkhunny):

There are two. The fundamental premise of a quadratic equation is that it has TWO solutions. Just because they happen to be equal, does not mean there is only one.

OpenStudy (anonymous):

A:121 B:121/4 c:-11/2 d:121/2 and halk of 11 is 5.5

OpenStudy (anonymous):

oh 11/2

OpenStudy (misty1212):

rigth

OpenStudy (misty1212):

now what is \(\left(\frac{11}{2}\right)^2\)?

OpenStudy (anonymous):

@tkhunny you mean the first one? and @misty1212 30.25

OpenStudy (misty1212):

lol stop with the decimals

OpenStudy (misty1212):

you see any decimals in your answer? no, of course not square the top, square the bottom

OpenStudy (anonymous):

30 1/4

OpenStudy (misty1212):

grrr

OpenStudy (misty1212):

fractions, not decimals not mixed numbers just fractions

OpenStudy (misty1212):

it is actually less work to do it with fractions \[\frac{11}{2}\times \frac{11}{2}=?\]

OpenStudy (anonymous):

121/4

OpenStudy (misty1212):

bingo

OpenStudy (anonymous):

thats the answer? cool, that was easy

OpenStudy (anonymous):

even though its a negative 11?

OpenStudy (misty1212):

lol yes it certainly was wan't it? when you know what you are doing much of this is very easy \[\color\magenta\heartsuit\]

OpenStudy (misty1212):

when you square it, it is positive

OpenStudy (anonymous):

cool thanks :)

OpenStudy (misty1212):

yw dear

OpenStudy (anonymous):

@misty1212 can you help me again?

OpenStudy (misty1212):

yes, but post in a new thread it is hard to scroll down all the time